### v1

# Keep the beat going: Automatic drum transcription with momentum

Abstract
A simple, interpretable way to perform automatic drum transcription is by factoring the magnitude spectrogram of a recorded musical piece using a partially fixed nonnegative matrix factorization. There are two natural ways to optimize the nonnegative matrix factorization, including a multiplicative update rule and projected gradient descent with momentum. The methods differ in their empirical accuracies and theoretical convergence guarantees. This paper summarizes the methods and their time complexities, and it applies the methods to the ENST-Drums data set and an original recording from the author’s band, evaluating the empirical accuracy with respect to ground-truth drum annotations. The results indicate that projected gradient descent with momentum leads to higher accuracy for a fixed runtime, and it satisfies stronger convergence guarantees.
1 Introduction
Nonnegative matrix factorization (NMF) is a data science approach that approximates a matrix as a product of two nonnegative matrices. NMF was introduced in 1994 by Paatero and Tapper as “positive matrix factorization” [14], and it was popularized by Lee and Seung [9] in 2000. NMF has many applications in music analysis [4], image representation [12], recommender systems [13], and genomics [18].
This work will focus on applying NMF to automatic drum transcription, a subfield within automatic music transcription [16]. An automatic drum transcription system could be useful to drummers as a tool for notating their grooves, as well as learning beats from recorded musical pieces. Previous work [15, 20, 19] has applied NMF to drum source detection, separation, and transcription. Researchers have also extended the NMF approach to automatic drum transcription using support vector machines [7] and convolutional neural networks [8].
This paper will investigate the partially fixed NMF method for automatic drum transcription developed in [20, 19]. There are two natural strategies to optimize the partially fixed NMF: (1) a multiplicative update rule based on gradient descent and (2) a projected gradient descent approach with Nesterov momentum. The main goal of this work is to compare the convergence of the two optimization algorithms. The results show that projected gradient descent with Nesterov momentum has favorable convergence guarantees and outperforms the multiplicative update rule in its empirical accuracy with a fixed runtime.
The paper is organized as follows. Section 2 describes the problem formulation, Section 3 summarizes optimization methods for NMF, Section 4 applies the methods to automatic drum transcription of the ENST-Drums database [5] and a demo of “Every Now and Then” by the author’s band, Threadbare [17], and Section 5 offers concluding remarks.
1.1 Notation
Scalars are in regular typeface: $n,d,L$. Vectors are in bold lower-case letters: $\bm{x},\bm{y}$. Matrices are in bold capital letters: $\bm{M},\bm{N}$. The $i$th entry of a vector $\bm{x}$ is $\bm{x}(i)$. The $(i,j)$ entry of a matrix $\bm{M}$ is $\bm{M}(i,j)$, the $i$th row is $\bm{M}(i,\cdot)$, and the $j$th column is $\bm{M}(\cdot,j)$. The transpose is $\bm{M}^{\top}$, the Moore-Penrose pseudoinverse is $\bm{M}^{+}$, the spectral norm is $\lVert\bm{M}\rVert$, and the Frobenius norm is $\lVert\bm{M}\rVert_{\rm F}$ or $\lVert\bm{x}\rVert_{2}$. The space of nonnegative-valued $m\times n$ matrices is denoted +$\mathbb{R}^{m\times n}_{+}$. The element-wise product of two vectors $\bm{x},\bm{y}$ is $\bm{x}\odot\bm{y}$. The element-wise product of two matrices $\bm{M},\bm{N}$ is $\bm{M}\odot\bm{N}$, and the element-wise division is $\frac{\bm{M}}{\bm{N}}$.
2 Problem formulation
The classic NMF problem is
(1) $\min_{\bm{W},\bm{H}}\lVert\bm{V}-\bm{WH}\rVert_{\rm F}^{2}$
where $\bm{V}$ is an $m\times n$ matrix and $\bm{W}\in\mathbb{R}^{m\times r}$ and $\bm{H}\in\mathbb{R}^{r\times n}$ are constrained to have all nonnegative entries.
In the context of automatic music transcription, $\bm{V}$ is the magnitude spectrogram matrix of a piece of music that has $m$ frequency bins and $n$ time frames. $\bm{W}$ is an $m\times r$ dictionary matrix that contains a basis of frequencies for the song, and $\bm{H}$ is an $r\times n$ activation matrix containing the temporal onsets of each component of $\bm{W}$ [16]. In this context, it is natural that $\bm{V}$ is approximated by the product of $\bm{W}$ and $\bm{H}$ matrices that both have nonnegative entries.
In 2010, Yoo et al. [20] introduced a method to separate drums from the rest of a musical track by assigning certain columns of the dictionary matrix and certain rows of the activation matrix to the drums. In 2015, Wu and Lerch [19] applied a similar approach, where they defined dictionary and activation matrices
$\bm{W}=\begin{bmatrix}\bm{W}_{D}\quad\bm{W}_{H}\end{bmatrix}\quad\text{and}\quad\bm{H}=\begin{bmatrix}\bm{H}_{D}\\ \bm{H}_{H}\end{bmatrix}.$
Here $\bm{W}_{D}$ is an $m\times r_{D}$ matrix whose columns indicate the frequency bases of each part of the drum kit. Typically the rank $r_{D}$ is the number of drum parts (snare drum, bass drum, hi-hat, etc.) to be transcribed. $\bm{W}_{H}$ contains the frequency bases of the harmonic instruments in the song, and its rank, $r_{H}$, is tuned to balance accuracy with interpretability. $\bm{H}_{D}$ is an $r_{D}\times n$ matrix that indicates when the components of $\bm{W}_{D}$ are active. Similarly, $\bm{H}_{H}$ is the $r_{H}\times n$ activation matrix for the harmonic instruments of the song.
In their “partially fixed NMF” approach, Wu & Lerch [19] train the matrix $\bm{W}_{D}$ using the ENST drums data set [5], and they keep this matrix fixed as $\bm{H}_{D},\bm{W}_{H},\bm{H}_{H}$ are updated. Therefore, the problem of interest Eq. 1 is modified and reformulated as minimizing a loss function
(2) $F(\bm{H}_{D},\bm{W}_{H},\bm{H}_{H})=\tfrac{1}{2}\lVert\bm{V}-(\bm{W}_{D}\bm{H}_{D}+\bm{W}_{H}\bm{H}_{H})\rVert_{\rm F}^{2},$
subject to the constraints that $\bm{H}_{D}$, $\bm{W}_{H}$, $\bm{H}_{H}$ are all nonnegative-valued. Yoo et al. [20] also consider methods of adapting $\bm{W}_{D}$ as the algorithm iterates, but this study will only consider fixed $\bm{W}_{D}$.
The major benefit of the partially fixed NMF approach is that having a pre-trained $\bm{W}_{D}$ allows for greater interpretability of the components. For example, the first column might represent the snare drum, the second column might represent the bass drum, etc. This ordering is necessary for transforming the data into human-readable drum notation. In practice, a drummer could calibrate $\bm{W}_{D}$ in the automatic drum transcription system by recording some drum hits to create a frequency basis for each part of a drum kit.
3 Summary of optimization methods
It is extremely challenging to find the global optimizer of the partially fixed NMF problem Eq. 2, since the loss function is nonconvex. Nevertheless, researchers have proposed finding a local optimizer using an alternating minimization [20]:
(3) $\bm{H}_{D}^{t+1}$ $=\operatorname*{arg\,min}_{\bm{H}_{D}\geq 0}F(\bm{H}_{D},\bm{W}_{H}^{t},\bm{H}_{H}^{t})$
$\bm{W}_{H}^{t+1}$ $=\operatorname*{arg\,min}_{\bm{W}_{H}\geq 0}F(\bm{H}_{D}^{t+1},\bm{W}_{H},\bm{H}_{H}^{t})$
$\bm{H}_{H}^{t+1}$ $=\operatorname*{arg\,min}_{\bm{H}_{H}\geq 0}F(\bm{H}_{D}^{t+1},\bm{W}_{H}^{t+1},\bm{H}_{H})$
Each subproblem in Eq. 3 is convex, so it can be solved relatively efficiently. There are two natural strategies to solve each subproblem based on a multiplicative update rule (MUR) or projected gradient descent with Nesterov momentum (NeNMF). The rest of this section will derive the MUR and NeNMF methods in detail and derive the best available mathematical convergence guarantees, building on techniques in [9, 6].
3.1 Multiplicative update rule
Lee and Seung [9] studied the classic NMF problem Eq. 1, and they proved the following result.
Theorem 3.1 (Convergence of original MUR algorithm [9]).
The square Frobenius norm $\|\bm{V}-\bm{WH}\|^{2}$ is nonincreasing under the multiplicative update rule
(4) $\bm{H}\leftarrow\bm{H}\odot\frac{\bm{W}^{\top}\bm{V}}{\bm{W}^{\top}\bm{WH}},\qquad\bm{W}\leftarrow\bm{W}\odot\frac{\bm{VH}^{\top}}{\bm{WHH}^{\top}}.$
In a critique of the MUR method, Lin [11] pointed out that a denominator term might equal zero, in which case the method is ill-defined. However, a zero denominator only arises in pathological settings, and Theorem 3.1 remains valid as long as the denominators stay positive.
Algorithm 1 MUR for partially fixed NMF
1:  Construct $\bm{W}_{D}\in\mathbb{R}_{+}^{m\times r_{D}}$, where $r_{D}$ is the number of drum components
2:  Randomly initialize $\bm{H}_{D}^{1}\in\mathbb{R}_{+}^{r_{D}\times n},\bm{W}_{H}^{1}\in\mathbb{R}^{m\times r_{H}}_{+},\bm{H}_{H}^{1}\in\mathbb{R}^{r_{H}\times n}_{+}$
3:  for $t=1,...,T$ do
4:     $\bm{H}_{D}^{t+1}=\bm{H}_{D}\odot\frac{\bm{W}_{D}^{\top}\bm{V}}{\bm{W}_{D}^{\top}(\bm{W}_{D}\bm{H}_{D}^{t}+\bm{W}_{H}^{t}\bm{H}_{H}^{t})}$
5:     $\bm{W}_{H}^{t+1}=\bm{W}_{H}^{t}\odot\frac{\bm{V}(\bm{H}_{H}^{t})^{\top}}{(\bm{W}_{D}\bm{H}_{D}^{t+1}+\bm{W}_{H}^{t}\bm{H}_{H}^{t}){(\bm{H}_{H}^{t})^{\top}}}$
6:     $\bm{H}_{H}^{t+1}=\bm{H}_{H}^{t}\odot\frac{({\bm{W}_{H}^{t+1}})^{\top}\bm{V}}{({\bm{W}_{H}^{t+1}})^{\top}(\bm{W}_{D}\bm{H}_{D}^{t+1}+\bm{W}_{H}^{t+1}\bm{H}_{H}^{t})}$
7:  end for
8:  $\bm{H}_{D}=\bm{H}_{D}^{T}$, $\bm{W}_{H}=\bm{W}_{H}^{T}$, $\bm{H}_{H}=\bm{H}_{H}^{T}$.
9:  return $\bm{H}_{D},\bm{W}_{H},\bm{H}_{H}$
In 2010, Yoo et al. [20] proposed solving the partially fixed NMF problem Eq. 2 using a variant of the multiplicative update rule Eq. 4. See Algorithm 1 for pseudocode derived from their work. As a limitation, however, the paper [20] did not provide a complete convergence analysis. To fill in this gap, the current paper proves a theoretical guarantee for MUR method by elaborating on the proof strategy in [9, Thm. 1].
Theorem 3.2 (Convergence of MUR algorithm for partially fixed NMF).
The loss function $F(\bm{H}_{D},\bm{W}_{H},\bm{H}_{H})=\frac{1}{2}\|\bm{V}-(\bm{W}_{D}\bm{H}_{D}+\bm{W}_{H}\bm{H}_{H})\|^{2}_{F}$ is nonincreasing under the multiplicative update rule
$\bm{H}_{D}$ $\leftarrow\bm{H}_{D}\odot\frac{\bm{W}_{D}^{\top}\bm{V}}{\bm{W}_{D}^{\top}(\bm{W}_{D}\bm{H}_{D}+\bm{W}_{H}\bm{H}_{H})},$
$\bm{W}_{H}$ $\leftarrow\bm{W}_{H}\odot\frac{\bm{VH}_{H}^{\top}}{(\bm{W}_{D}\bm{H}_{D}+\bm{W}_{H}\bm{H}_{H})\bm{H}_{H}^{\top}},$
$\bm{H}_{H}$ $\leftarrow\bm{H}_{H}\odot\frac{\bm{W}_{H}^{\top}\bm{V}}{\bm{W}_{H}^{\top}(\bm{W}_{D}\bm{H}_{D}+\bm{W}_{H}\bm{H}_{H})},$
as long as the denominator terms are positive.
Proof 3.3.
We will prove that the update rule applied to $\bm{H}_{D}$ does not increase the loss function $F(\bm{H}_{D},\bm{W}_{H},\bm{H}_{H})$. A similar proof shows that the update rule applied to $\bm{W}_{H}$ or $\bm{H}_{H}$ does not increase the loss function.
We observe that the loss function is separable in the columns of $\bm{H}_{D}$:
$F(\bm{H}_{D},\bm{W}_{H},\bm{H}_{H})=\frac{1}{2}\sum_{j=1}^{n}\|\bm{V}(\cdot,j)-[\bm{W}_{D}\bm{H}_{D}(\cdot,j)+\bm{W}_{H}\bm{H}_{H}(\cdot,j)]\|_{2}^{2}.$
We consider the change to the loss function after updating a single column from $\bm{H}_{D}(\cdot,j)$ to $\bm{x}\in\mathbb{R}^{r}$:
$\tilde{F}(\bm{x})=F(\bm{H}_{D}+[\bm{x}-\bm{H}_{D}(\cdot,j)]\bm{e}_{j}^{\top},\bm{W}_{H},\bm{H}_{H})$
$\tilde{F}(\bm{x})$ is a quadratic function of $\bm{x}$, and the gradient and Hessian terms are given as follows:
$\nabla\tilde{F}(\bm{x})=\bm{W}_{D}^{\top}[\bm{W}_{D}\ \bm{x}+\bm{W}_{H}\bm{H}_{H}(\cdot,j)-\bm{V}(\cdot,j)],\qquad\nabla^{2}\tilde{F}(\bm{x})=\bm{W}_{D}^{\top}\bm{W}_{D}.$
Hence, $\tilde{F}(\bm{x})$ is specified by the formula
$\tilde{F}(\bm{x})$ $=\tilde{F}(\bm{H}_{D}(\cdot,j))+[\bm{x}-\bm{H}_{D}(\cdot,j)]^{\top}\nabla\tilde{F}(\bm{H}_{D}(\cdot,j))$
$+\frac{1}{2}[\bm{x}-\bm{H}_{D}(\cdot,j)]^{\top}\bm{W}_{D}^{\top}\bm{W}_{D}[\bm{x}-\bm{H}_{D}(\cdot,j)].$
To complete the proof, it suffices to show that updating the $j$th column from $\bm{H}_{D}(\cdot,j)$ to
(5) $\bm{x}_{\star}=\bm{H}_{D}(\cdot,j)\odot\frac{\bm{W}_{D}^{\top}\bm{V}(\cdot,j)}{\bm{W}_{D}^{\top}[\bm{W}_{D}\bm{H}_{D}(\cdot,j)+\bm{W}_{H}\bm{H}_{H}(\cdot,j)]}$
leads to a loss function $\tilde{F}(\bm{x}_{\star})\leq\tilde{F}(\bm{H}_{D}(\cdot,j))$.
We introduce a diagonal matrix $\bm{K}\in\mathbb{R}^{r\times r}$ with elements
$\bm{K}(i,i)=\begin{cases}\frac{\bm{W}_{D}(\cdot,i)^{\top}[\bm{W}_{D}\bm{H}_{D}(\cdot,j)+\bm{W}_{H}\bm{H}_{H}(\cdot,j)]}{\bm{H}_{D}(i,j)},&\bm{H}_{D}(i,j)>0,\\ 0,&\bm{H}_{D}(i,j)=0.\end{cases}$
For any multiplicative update $\bm{x}=\bm{H}_{D}(\cdot,j)\odot\bm{y}$ with $\bm{y}\in\mathbb{R}^{r}$, we can upper bound
$(\bm{H}_{D}(\cdot,j)\odot\bm{y})^{\top}\bm{W}_{D}^{\top}\bm{W}_{D}(\bm{H}_{D}(\cdot,j)\odot\bm{y})$
$=\sum_{i=1}^{r}\sum_{\ell=1}^{r}\bm{y}(i)\bm{y}(\ell)\bm{H}_{D}(i,j)\bm{W}_{D}(\cdot,i)^{\top}\bm{W}_{D}(\cdot,\ell)\bm{H}_{D}(\ell,j)$
$\leq\sum_{i=1}^{r}\sum_{\ell=1}^{r}\biggl{[}\frac{1}{2}\bm{y}(i)^{2}+\frac{1}{2}\bm{y}(\ell)^{2}\biggr{]}\bm{H}_{D}(i,j)\bm{W}_{D}(\cdot,i)^{\top}\bm{W}_{D}(\cdot,\ell)\bm{H}_{D}(\ell,j)$
$=\sum_{i=1}^{r}\sum_{\ell=1}^{r}\bm{y}(i)^{2}\bm{H}_{D}(i,j)\bm{W}_{D}(\cdot,i)^{\top}\bm{W}_{D}(\cdot,\ell)\bm{H}_{D}(\ell,j)$
$\leq\sum_{i=1}^{r}\sum_{\ell=1}^{r}\bm{y}(i)^{2}\bm{H}_{D}(i,j)\bm{W}_{D}(\cdot,i)^{\top}[\bm{W}_{D}(\cdot,\ell)\bm{H}_{D}(\ell,j)+\bm{W}_{H}(\cdot,\ell)\bm{H}_{H}(\ell,j)]$
$=(\bm{H}_{D}(\cdot,j)\odot\bm{y})^{\top}\bm{K}(\bm{H}_{D}(\cdot,j)\odot\bm{y}).$
Therefore, for any multiplicative update $\bm{x}=\bm{H}_{D}(\cdot,j)\odot\bm{y}$, the quadratic function
$G(\bm{x})=\tilde{F}(\bm{H}_{D}(\cdot,j))+[\bm{x}-\bm{H}_{D}(\cdot,j)]^{\top}\nabla\tilde{F}(\bm{H}_{D}(\cdot,j))+\frac{1}{2}[\bm{x}-\bm{H}_{D}(\cdot,j)]^{\top}\bm{K}[\bm{x}-\bm{H}_{D}(\cdot,j)]$
provides an upper bound on the loss function $\tilde{F}(\bm{x})$.
Last, the minimizer of $\bm{y}\mapsto G(\bm{H}_{D}(\cdot,j)\odot\bm{y})$ satisfies the gradient equal to zero condition:
$0=\bm{H}_{D}(\cdot,j)\odot\bm{K}[\bm{K}^{+}\nabla\tilde{F}(\bm{H}_{D}(\cdot,j)+\bm{H}_{D}(\cdot,j)\odot\bm{y}-\bm{H}_{D}(\cdot,j)].$
Equivalently, the minimizer satisfies the following identity:
$\bm{H}_{D}(\cdot,j)\odot\bm{y}$ $=\bm{H}_{D}(\cdot,j)-\bm{K}^{+}\nabla\tilde{F}(\bm{H}_{D}(\cdot,j)$
$=\bm{H}_{D}(\cdot,j)-\frac{\bm{H}_{D}(\cdot,j)\odot\bm{W}_{D}^{\top}[\bm{W}_{D}\bm{H}_{D}(\cdot,j)+\bm{W}_{H}\bm{H}_{H}(\cdot,j)-\bm{V}(\cdot,j)]}{\bm{W}_{D}^{\top}[\bm{W}_{D}\bm{H}_{D}(\cdot,j)+\bm{W}_{H}\bm{H}_{H}(\cdot,j)]}$
$=\bm{H}_{D}(\cdot,j)\odot\frac{\bm{W}_{D}^{\top}\bm{V}(\cdot,j)}{\bm{W}_{D}^{\top}[\bm{W}_{D}\bm{H}_{D}(\cdot,j)+\bm{W}_{H}\bm{H}_{H}(\cdot,j)]}.$
Hence, the minimizer of $G$ is precisely the multiplicative update rule $\bm{x}_{\star}=\bm{H}_{D}(\cdot,j)\odot\bm{y}$ given in Eq. 5. Since $G$ is an upper bound on $\tilde{F}$ which intersects at $\bm{H}_{D}(\cdot,j)$, we conclude:
$\tilde{F}(\bm{x}_{\star})\leq G(\bm{x}_{\star})\leq G(\bm{H}_{D}(\cdot,j))=\tilde{F}(\bm{H}_{D}(\cdot,j)).$
This shows that the multiplicative update rule does not increase the loss function.
3.2 Projected gradient descent with momentum
Algorithm 2 $\operatorname{OGM}(\bm{W},\bm{H}_{0},\bm{V})$
1:  Initialize $\alpha_{1}=1$, $\bm{Y}_{1}=\bm{H}_{0}$
2:  for $k=1,...,K$ do
3:     $\bm{H}_{k}=\operatorname{ReLu}\Bigl{(}\bm{Y}_{k}-\frac{\bm{W}^{\top}[\bm{W}\bm{Y}_{k}-\mathbf{V}]}{\lVert\bm{W}\rVert^{2}}\Bigr{)}$
4:     $\alpha_{k+1}=\frac{1+\sqrt{4\alpha_{k}^{2}+1}}{2}$
5:     $\bm{Y}_{k+1}=\bm{H}_{k}+\frac{\alpha_{k}-1}{\alpha_{k+1}}(\bm{H}_{k}-\bm{H}_{k-1})$
6:  end for
7:  $\bm{H}=\bm{H}_{K}$
8:  return $\bm{H}$
In 2012, Guan and coauthors [6] introduced the “NeNMF” method for the classic NMF problem Eq. 1. Like MUR, the NeNMF method performs an alternating minimization over $\bm{W}$ and $\bm{H}$ factors. As the major innovation, NeNMF solves each subproblem
$\min_{\bm{H}\geq 0}\lVert\bm{W}\bm{H}-\bm{V}\rVert_{\rm F}^{2}\quad\text{or}\quad\min_{\bm{W}\geq 0}\lVert\bm{W}\bm{H}-\bm{V}\rVert_{\rm F}^{2}$
to high accuracy. To do this, NeNMF applies several iterations of proximal gradient descent with Nesterov momentum, which Guan et al. call the “optimal gradient method” (OGM).
Algorithm 2 provides pseudocode for OGM. The method is based on gradient descent, and it projects each iterate onto the cone of nonnegative-valued matrices by an element-wise application of the rectified linear unit $\operatorname{ReLu}(x)=\max\{x,0\}$. The method incorporates Nesterov momentum using a step size parameter $\alpha_{k}$, which promotes an accelerated convergence rate. Guan et al. established the following guarantee for OGM in [6, Prop. 2.1].
Theorem 3.4 (Convergence of OGM [6]).
Let $\bm{W}\in\mathbb{R}^{m\times r}$, $\bm{H}_{0}\in\mathbb{R}_{+}^{r\times n}$, and $\bm{V}\in\mathbb{R}^{m\times n}$ be inputs to the optimal gradient method (Algorithm 2), and fix a matrix
$\bm{H}_{\star}\in\operatorname*{arg\,min}_{\bm{H}\geq\bm{0}}\lVert\bm{W}\bm{H}-\bm{V}\rVert_{\rm F}^{2}.$
Then, after $K$ iterations, the optimal gradient method outputs a matrix $\bm{H}=\bm{H}_{K}$ that satisfies
$\lVert\bm{W}\bm{H}_{K}-\bm{V}\rVert_{\rm F}^{2}\leq\lVert\bm{W}\bm{H}_{\star}-\bm{V}\rVert_{\rm F}^{2}+\frac{2\lVert\bm{W}\rVert^{2}\lVert\bm{H}_{0}-\bm{H}_{\star}\rVert_{\rm F}^{2}}{(K+2)^{2}}.$
Theorem 3.4 ensures that OGM solves each subproblem up to an additive error of $\mathcal{O}\big{(}\frac{1}{K^{2}}\big{)}$ after $K$ iterations.
To our knowledge, NeNMF has not previously been applied to the partially fixed NMF problem Eq. 2. However, we can extend NeNMF to the partially fixed NMF problem by using OGM to solve each subproblem in the alternating minimization Eq. 3. Algorithm 3 presents pseudocode for this new NeNMF extension. Because Algorithm 3 is using OGM to solve each subproblem, convergence guarantees follow immediately from Theorem 3.4.
Algorithm 3 NeNMF for partially fixed NMF
1:  Construct $\bm{W}_{D}\in\mathbb{R}_{+}^{m\times r_{D}}$, where $r_{D}$ is the number of drum components
2:  Randomly initialize $\bm{H}_{D}^{1}\in\mathbb{R}_{+}^{r_{D}\times n},\bm{W}_{H}^{1}\in\mathbb{R}^{m\times r_{H}}_{+},\bm{H}_{H}^{1}\in\mathbb{R}^{r_{H}\times n}_{+}$
3:  for $t=1,...,T$ do
4:     $\bm{H}_{D}^{t+1}=\operatorname{OGM}(\bm{W}_{D},\bm{H}_{D}^{t},\bm{V}-\bm{W}_{H}^{t}\bm{H}_{H}^{t})$
5:     $\bm{W}_{H}^{t+1}=\operatorname{OGM}((\bm{H}_{H}^{t})^{\top},(\bm{W}_{H}^{t})^{\top},(\bm{V}-\bm{W}_{D}\bm{H}_{D}^{t+1})^{\top})^{\top}$
6:     $\bm{H}_{H}^{t+1}=\operatorname{OGM}(\bm{W}_{H}^{t+1},\bm{H}_{H}^{t},\bm{V}-\bm{W}_{D}\bm{H}_{D}^{t+1})$
7:  end for
8:  $\bm{H}_{D}=\bm{H}_{D}^{T}$, $\bm{W}_{H}=\bm{W}_{H}^{T}$, $\bm{H}_{H}=\bm{H}_{H}^{T}$.
9:  return $\bm{H}_{D},\bm{W}_{H},\bm{H}_{H}$
3.3 Comparison of runtimes and convergence
Updating a single matrix in an iteration of MUR or OGM requires $\mathcal{O}(mn)$ arithmetic operations, and the updates have similar empirical runtimes. Since NeNMF performs $K$ iterations of OGM, the cost of NeNMF is roughly $K$ times higher than the cost of MUR. The results in Section 4 account for this factor-of-$K$ difference by running MUR for $T=100$ iterations versus running NeNMF with $T=10$ outer iterations containing $K=10$ inner OGM iterations.
Obtaining the global minimizer of the partially fixed NMF problem Eq. 2 is challenging since the loss function is nonconvex. Nevertheless, we can hope to solve each convex subproblem in the alternating minimization Eq. 3 to high accuracy. NeNMF is guaranteed to solve each subproblem up to an additive error of $\mathcal{O}(1/K^{2})$ due to Theorem 3.4. In contrast, MUR is not guaranteed to solve each subproblem to any level of accuracy, although it does decrease the overall loss due to Theorem 3.2.
4 Application to automatic drum transcription
This section describes numerical experiments applying the partially fixed NMF for automatic drum transcription. Section 4.1 introduces data sets and Section 4.2 presents results.
4.1 Data sets
ENST-Drums [5] is a publicly available data set containing annotated audio recordings from three drummers playing individual hits, short phrases, and longer songs in various styles on their own drum kits. We downloaded the minus-one subset of the ENST-Drums data from the repository [1]. This subset contains 28 tracks of length $\approx 70$ s with the drummers playing their drum kits without accompaniment from other instruments. The sample rate is 16 kHz.
“Every Now and Then” is a song from the author’s band, Threadbare [17]. It consists of two guitars, one bass guitar, vocals, and a drum kit, and it is available for streaming at https://open.spotify.com/track/3FoWRPLO9Z0tQiVY9jLrth?si=1e252b92cde04a8e. The sample rate is 44.1 kHz. We manually annotated each onset of the snare drum or bass drum. See Fig. 1 for a labeled photo of the complete drum kit, including the two annotated components.
Figure 1: Photograph of the author’s drum kit with instruments annotated by hand. The snare and bass drums were included in the automatic drum transcription for the song “Every Now and Then”, as described in Section 4.
As a preprocessing procedure, we converted ENST-Drums and “Every Now and Then” into magnitude spectrograms using the short-time Fourier transform with a window length 2048 and hop size 512. This transformation resulted in a time resolution of 32 ms for ENST-Drums and 12 ms for “Every Now and Then”.
Last, we constructed a dictionary element for each annotated drum component by extracting the magnitude spectrogram of an individual hit and averaging over the time axis. We constructed dictionary elements for the hi-hat, snare drum, and bass drum for each of the three drum kits in ENST-Drums. We constructed dictionary elements for the snare drum and bass drum for “Every Now and Then”.
4.2 Results
Figure 2: Square Frobenius norm error as a function of runtime for MUR and NeNMF. The horizontal axis shows the number of iterations for MUR and 10×10$\times$10 × the number of outer iterations for NeNMF. The error for ENST-Drums (left) is averaged across 28 tracks, while the error for “Every Now and Then” (right) is for a single track.
We applied the MUR and NeNMF algorithms to solve the partially fixed NMF problem Eq. 2 for ENST-Drums and “Every Now and Then” data using a harmonic rank $r_{H}=5$. In these tests, we independently initialized the entries of $\bm{H}_{D},\bm{W}_{H},$ and $\bm{H}_{H}$ as $\operatorname{Unif}(0,1)$ random variables. Then we applied either MUR for 100 iterations or NeNMF for 10 outer iterations using 10 inner OGM iterations. These experiments led to similar runtimes for the two algorithms.
Figure 2 compares the square Frobenius norm errors achieved using MUR and NeNMF. The errors for ENST-Drums (left panel) are averaged across the 28 tracks, while the errors for “Every Now and Then” (right panel) are for a single track. To promote a fair comparison, the horizontal axis records the iteration parameter $t$ in MUR (Algorithm 1) and it records 10×10$\times$10 × the iteration parameter $t$ in NeNMF (Algorithm 3). The results show that NeNMF leads to a comparable or smaller square Frobenius norm error for either data set and any fixed runtime.
Next, we evaluated the accuracy of the drum hits identified using MUR and NeNMF. Both optimization algorithms produce drum onset matrices consisting of nonnegative entries $\bm{H}_{D}(i,j)$. As a complication, however, there are various ways to determine which entries are large enough to be regarded as drum hits or “onsets”. One standard method used in automatic music transcription [3, 2] identifies $p_{j}$ of the largest entries in each column $\bm{H}_{D}(\cdot,j)$, where $p_{j}$ is the ground-truth number of drum hits. The ground-truth number of drum hits is not likely to be available in real-world use cases. Yet this approach allows for straightforward comparisons without the need to tune evaluation hyperparameters. After identifying the drum onsets using the ground-truth number of onsets, we counted the true positives (TP), false positives (FP), and false negatives (FN) in each row of the onset matrix $\bm{H}_{D}$. Then, we measured accuracy using the $F$-score:
$F=\frac{2{\rm TP}}{2{\rm TP}+{\rm FP}+{\rm FN}}.$
Each $F$-score lies in the range $F\in[0,1]$, with higher $F$-scores indicating greater accuracies.
Table 1 reports the F-scores for each data set and optimization method, averaged over all the drum components and tracks for each data set. The results show that NeNMF achieves higher accuracy than MUR for both ENST-Drums and “Every Now and Then”. NeNMF achieves an especially high accuracy of 97.5% for “Every Now and Then”, because the track contains just two annotated drum components and the onset matrix correctly indicates which component is larger at nearly all the onset times.
MUR NeNMF
ENST-Drums 0.599 0.620
“Every Now and Then” 0.854 0.975
Table 1: F-scores evaluated using ground-truth number of offsets. Higher F-scores are marked in bold.
A different way to identify drum hits is based on a signal-adaptive median threshold [19, 10]. This method evaluates the median in a sliding window over each row $\bm{H}_{D}(i,\cdot)$, and it adds a small offset to determine the threshold for a drum hit. For the experiments, we set the window length to 0.1 s, and we set the offset coefficients to 0.05 for the hi-hat, 0.1 for the snare drum, and 0.15 for the bass drum. We considered a detected onset correct if it is within 50 ms of a ground-truth onset. Note that changing the window length and the offset coefficients can lead to large changes in the evaluation scores, and different tracks have better scores for different values.
MUR NeNMF
Hi-hat 0.254 0.283
Snare drum 0.308 0.298
Bass drum 0.290 0.629
Table 2: F-scores for ENST-Drums using the signal-adaptive median threshold evaluation. Higher F-scores are marked in bold.
MUR NeNMF
Snare drum 0.275 0.369
Bass drum 0.444 0.454
Table 3: F-scores for “Every Now and Then” using the signal-adaptive median threshold evaluation. Higher F-scores are marked in bold.
Tables 3 and 3 present the F-scores evaluated using the signal-adaptive median threshold. The scores are averaged over 28 tracks for ENST-Drums. The F-scores are generally higher for NeNMF than for MUR. The highest F-score of 62.9% is achieved using NeNMF to detect the bass drum in ENST-Drums, and it is over 2×2$\times$2 × higher than the corresponding F-score of 29.0% achieved using MUR.
5 Conclusion
In our empirical results, NeNMF is generally more accurate than MUR given a fixed computational budget. Further, NeNMF offers advantageous convergence guarantees. Therefore, we recommend using NeNMF for automatic drum transcription and other partially fixed NMF applications in the future.
Acknowledgments
We would like to acknowledge Dan Manil, Zachary Lawrence, and Shawn Pana for collaboration in writing, performing, and recording “Every Now and Then.”